OpenStudy (anonymous):

Determine the domain of the function. f(x)=(x-5)/(^4√x^2-9). I need it in interval notation. I got (-∞,-3)(3,∞) but I'm not sure it's right.

OpenStudy (anonymous):

\[f(x)=\frac{ x-5 }{ \sqrt[4]{x ^{2}-9} }\]

OpenStudy (abb0t):

Your domain is all numbers which make your function undefined. And anything divided by zero would make that true. So yes, \(\pm 3\) would be your domain from \(-\infty ~ to~ \infty\)

OpenStudy (anonymous):

Thank you very much!:)

OpenStudy (preetha):

Dont forget to give a medal!

OpenStudy (anonymous):

Just a minor correction, @abb0t, the domain is all numbers that make the function *not* undefined.